24cosA +7sin B what is value of this question please give explanation.i am grateful of you

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Any expression of the form
a sin x + bsinx
 
Its maximum value = sqrt(a²+b²)
Hence its minimum value = - sqrt(a² +b²)
 
Hence in the above question
Maximum value= sqrt(24² +7²) = sqrt(625)  = 25
 
Regards
Arun (askIITians forum expert)